You are here: Foswiki>Commissioning Web>Logs>22Oct2012_Log (23 Oct 2012, JuanGuerra)Edit Attach

UT 20121022 AO#1 Wakeup

Observer:JCGuerra (LBTO), A.Puglisi remote(INAF)
Telescope Operator: David G.(LBTO)
Additional AO Support: JCGuerra (LBTO)
Telescope Support: Michael.K (LBTO)
Instrument Support: John M (LBTO)

Summary

  • Retro On. Some problem to find the RR brackets.

  • Startup of the DX size.

  • Telescope setup to work with the RR.

  • Alfio on line.

  • The W unit is set up without any problem to load the board setup.

  • Afio update all the routines that we have to work with the RR

  • Right spot center.

  • Test of close loop, autofocus of IRTC. Results shows a 66% in H with a non-optimized loop. A more detailed optimization is planned for tomorrow.

  • We had some error with the TCS subsystems, telemetry was lost and the the primary went to panic.

  • Handover 3:30

Details

Start 08:00

  • Installation of the RR has been delayed till 9:30 due to the missing RR brackets. They were not in the usual storage location. This location is on the RR trolley.

  • The DX is power on the system came up without any problem.

  • W and telescope ready, the spot can not be seeing using the old hexapod position. The hexa_spiral procedure to acquire the spot is used. The procedure didn't work, is not installed properly in the DX side. Alfio is fixing this bug.

  • Spots found with hexapod position: 0.1,0,3.4,-303.9,-32,0. Hexapod position for the wide field of view in IRTC. Center the left spot. For the narrow filed of view the hexapod position are:
    -0.9,1.3,3.44,-336.7,-139.7,0 
    Center the right spot. For the narrow filed of view the hexapod position are:
    *-0.91,1.46,3.48,-363,-137.6.0* 

The pupil are digital shifted by 2 pixels and the rotaion offset changed by 1 degree. This is test with the Bin1, a more detailed check has to be done with the others bin. The new values are updated in teh w configuration files. Alfio is working on. (19:20)

The system close the loop with the a test configuration, there isn't any problem all the commands went through the system adn we could close the loop. The next item is to run the autofocus routine to improve the irtc image quality and to get some performance.

Center the left spot in the narrow field. Closing the loop and taken a few data, we can see the cross effect in the psf which is a high indicator that the spot is wrong. Data taken

20121022_202408 ??? 6.8 90 4 NaN -1.00 092604 1 400 990 0.4 0.1 -1.0 3 1182 ON 68.1 H 30 100 OFF NO - Pupils mismatch - Warning: Selected IRTC dark +16717h old!
20121022_202428 ??? 6.8 90 4 NaN -1.00 092604 1 400 990 0.4 0.1 -1.0 3 1182 ON 68.1 H 30 100 OFF NO - Pupils mismatch
WARNING: DISTURB IS OFF!!
20121022_202442 ??? 6.7 90 4 NaN -1.00 092604 1 400 990 0.4 0.1 -1.0 3 1183 ON 71.3 H 30 100 OFF NO - long stop of 80 frames at frame 3695 - Pupils mismatch

19:20 Primary went in panic.

21:00 Lost telescope information and rip of the shell. Some subsystem has to be reinizilaze.

The left spot is not the right one. The right spot is moved to the cebnter of teh narrow field and test of close loop. The focus of the psf is adjusted running the autofocus procedure. With the focus and the AO optimal parmaters we close the loop and we run autogain. The autogain works without any problem.

With a non-optimized system we close the loop and some data are taken. We are having problems with the acquisition routine. The errors are reported to Alfio and we expected to fix the bugs. A results of a non-optimized psf is showed in the table and images below:

Tracknum 20121022_223333
Closed Loop Yes
Is OK? No - Pupils mismatch
seeing [arcsec] 0.800000
Vwind [m/s] 14.840388
AntiDrift ON
nphotons/sub/fr 1251.624390
Magnitude 6.688661
# Modes 400
Modal rec Rec_20101024_092604.fits
Binning 1
Frequency [Hz] 989
Modulation 2.996182
FW1 Dichroic 600-1000 nm
FW2 empty
Gain minmax 0.000000 0.513000
Telescope elevation 90.004222
Wind speed NaN
lambda [um] 1.600000
exptime [s] 0.029999
framerate [Hz] 33.334446
no. frames 100
SR SE 0.661246
IRTC dark 20121022_223435.fits_cube.fits

* psf and modalplot:
Screenshot_testbin1b.png

handover 3:30

-- JuanGuerra - 22 Oct 2012
  • psf and modalplot:
    Screenshot_testbin1b.png
Topic revision: r5 - 23 Oct 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback